# Enhancing the Accuracy of Spatio-Temporal Models for Wind Speed Prediction by Incorporating Bias-Corrected Crowdsourced Data

### v1

@@ Appendix B Bias Correction Methods
1 $\lambda=\lambda^{GWA}$ 1.62
2 $\lambda\sim\beta_{0}+\beta_{1}\lambda^{GWA}$ $\hat{\lambda}=-3.1+1.28\lambda^{GWA}$ 1.04 65.5
3 $\lambda\sim\beta_{0}+\beta_{1}\lambda^{GWA}+\beta_{2}Dist$ $\hat{\lambda}=-0.93+1.03\lambda^{GWA}-0.03Dist$ 0.82 52.1
4 $\lambda\sim\beta_{0}+s\left(\lambda^{GWA}\right)$ $\hat{\lambda}=5.1+s(\lambda^{GWA})$ 0.78 54.7
5 $\lambda\sim\beta_{0}+s\left(\lambda^{GWA}\right)+\beta_{1}Dist$ $\hat{\lambda}=5.6+s(\lambda^{GWA})-0.02Dist$ 0.57 42.2
@@ Appendix F Modelling Data using Normal Transformation
1 Meteorological Stations only 1.38
2 Met and PWS (bias correction), single nugget variance 1.31
3 Met and PWS (bias correction), group specific variance 1.30
@@
Appendix G Performance of AR1 versus Independent Gaussian Processes
One of the notable observations in both the simulation study and the wind speed dataset is that the independent Gaussian process (IGP) model performs comparably or even outperforms the AR1 model, despite the latter’s theoretical advantage in capturing temporal correlation. In the simulation study, this result may be explained by the separable spatio-temporal process used for data generation. For a separable process, the value at an unobserved location conditionally depends only on other locations at the same time point, and its own past or future values at the same location. Since the test case involves predictions at a completely unobserved location, the temporal component of the AR1 model provides no additional information, reducing its predictive advantage over the purely spatial IGP model.
In the wind speed data, the estimated AR1 correlation parameter is very close to 1, suggesting the process is nearly a random walk. Notably, the spatial range parameter in the AR1 model is smaller than that of the IGP model, indicating that the AR1 model attributes most of the observed dependence to temporal persistence rather than spatial structure (See table 17). Consequently, predictions at unobserved locations often tend toward the station-specific mean, as the reduced spatial range limits borrowing of information across stations. This behaviour likely explains the IGP’s superior performance—its larger spatial range allows it to leverage cross-station correlations more effectively.
The limited range may also be due to how the AR1 parameters are inferred. The spatial process at time $t$ is expressed as
$z_{s,t}=\rho z_{s,t-1}+\xi_{s,t},$
$\xi_{s,t}$ represents a spatially correlated innovation term between $t$ and $t+1$, which is used to infer spatial hyperparameters. Studying the wind speed data, empirical analysis suggests the hourly innovations show low spatial range, despite stations showing the same long term trends. In Figure 19 the correlation of hourly changes against distance is compared in both the wind dataset and the simulated data from a true AR1 process, demonstrating how hourly changes in the wind dataset are less correlated than expected in simulated data.
Parameter Parameter Value - IGP Parameter Value - AR1
Effective Range (km) 456 115
$\sigma_{\text{z}}$ 0.325 0.85
$\sigma_{\text{Met}}$ 0.266 0.150
$\sigma_{\text{A}}$ 0.271 0.154
$\sigma_{\text{B}}$ 0.241 0.146
$\sigma_{\text{C}}$ 0.268 0.265
$\sigma_{\text{U}}$ 0.253 0.154
$\rho$ 0.98
Table 17: Estimated parameters for Model 6 under both the IGP and AR1 covariance structure.
(a) (a)
(b) (b)
Figure 19: A plot of empirical correlation against distance for the hourly increments for (a) the wind speed dataset and (b) the simulated data from a true AR1 process in Section 4.5.
Appendix H Parameter Confidence Intervals by Number of Timepoints

### v2

@@ Appendix B Bias Correction Methods
1 $\lambda=\lambda^{GWA}$ 1.07
2 $\lambda\sim\beta_{0}+\beta_{1}\lambda^{GWA}$ $\hat{\lambda}=-1.5+1.1\lambda^{GWA}$ 0.56 44.6
3 $\lambda\sim\beta_{0}+\beta_{1}\lambda^{GWA}+\beta_{2}Dist$ $\hat{\lambda}=-0.55+0.97\lambda^{GWA}-0.001Dist$ 0.53 44.2
4 $\lambda\sim\beta_{0}+s\left(\lambda^{GWA}\right)$ $\hat{\lambda}=5.1+s(\lambda^{GWA})$ 0.55 44.8
5 $\lambda\sim\beta_{0}+s\left(\lambda^{GWA}\right)+\beta_{1}Dist$ $\hat{\lambda}=5.3+s(\lambda^{GWA})-0.01Dist$ 0.49 42.2
@@ Appendix F Modelling Data using Normal Transformation
1 Meteorological Stations only 1.34
2 Met and PWS (bias correction), single nugget variance 1.28
3 Met and PWS (bias correction), group specific variance 1.27
@@
Appendix G Parameter Confidence Intervals by Number of Timepoints